Half a bottle of scotch whiskey is consumed and returned to the cabinet uncorked. Initially the scotch contained 400 grams of ethanol. The unsealed bottle loses ethanol at rate of 0.5grams per day. After sitting in the cabinet over the summer (90 days) with the cork removed, how much alcohol remains in the scotch?
Answer #1

Initial amount of alcohol = 400g

Amount of alcohol after consumption = 400/2 = 200g

Amount of alcohol lost = 0.5 * 90 = 45 g

Amount of alcohol remaining = 200-45 = 155g